I have on hand 98% alcohol. How much sterile water do I need to add to arrive at a solution of 20%.
Answer by ptaylor(2198)

Please re-check your problem.
Seems like we need to know the amount of 98% alcohol that you have on hand.
Hope this helps---ptaylor
The 98% alcohol is in a 5 ml vial, soooo
Let x=amount of sterile water needed
Now we know that the amount of pure alcohol before the mixture takes place(0.98)*5 has to equal the amount of pure alcohol after the mixture takes place 0.20(5+x).
Sooo our equation to solve is:
0.98*5=0.20(5+x) simplify
4.9=1+0.20x subtract 1 from each side
3.9=0.20x
x=19.5 ml---amount of sterile water needed
CK
0.98*5=0.20(5+19.5)
4.9=1+3.9
4.9=4.9
Hope this helps---ptaylor